An economy has two firms. Firm A produces wheat and firm B produces bread. In a given year, produces 50 bushels of wheat, sells 20 bushels to firm B at $3, exports 25 bushels at the same price to France, and stores 5 bushels as inventory. Firm A pays $50 in wages to consumers. Firm B produces 50 loaves of bread and sells them to consumer for $2 per loaf. Firm B pays $20 in wages. In addition, consumers import and consume 15 loaves from France paying $1 per loaf.

(d),Assume that the owners of Firm B are French and repatriate all of Firm B’s profits. What is the current account balance? What is the financial account balance?
Answer #1

Solution: Exports: Firm A exports 25 bushels * $3/bu. = $75

Imports: The import by consumers 5 loaves of bread * $1/loaf = $15.

Net exports = Exports - Imports = $75 - $15 = $60

Firm B pays $60 to Firm A for $20 intermediate input (bushels of wheat) thus the value added equals $40 (=$60 - $40)

Thus NFP = -$20

 

Current account = Net exports + Net factor payments = 60- 20 = $40

Financial account = - Current account = -$40